3.1. If R is a ring and a, b, c, d ∈ R, evaluate (a + b)(c + d).

3.2. Prove that if a, b ∈ R, then (a + b)2 = a2 + ab + ba + b2 where

by x2 we mean xx.


1
Expert's answer
2022-06-16T09:16:14-0400

3.1

"(a + b)(c + d) = a(c + d) + b(c + d)"

by distributive law:

"= (ac + ad) + (bc + bd) = ac + ad + bc + bd"


3.2

"(a + b)^2 = (a + b)(a + b) = a(a + b) + b(a + b)=a^2 + ab + ba + b^2"

Note that if R is not a commutative ring "ab\ne ba"